a beam, simply supported at the ends, carries a concentrated load of 300 lb/ft and a uniformly distributed load weighing 100 lb/ft as shown. determine the reactions at the beam
1
Expert's answer
2020-10-23T07:18:22-0400


Write the equilibrium of forces around the left support:


"R_L-300-100\cdot6+R_R=0."


Write the momentum equilibrium around the left support:


"0=300\cdot2+100\cdot6\cdot\bigg(2+2+\frac{6}2\bigg)-R_R\cdot10,\\\space\\
R_R=480\text{ lb},\\
R_L=420\text{ lb}."